An object is projected directly upward from te ground. After t secconds. the distance (s) from the ground is given by the equation: s=144t- 16t to the 2nd power. After how many seconds will the object reach 324 feet above the ground?
Answer by stanbon(48554)

An object is projected directly upward from the ground. After t secconds. the distance (s) from the ground is given by the equation: s=144t- 16t^2 . After how many seconds will the object reach 324 feet above the ground?
---------
s is the distance above the ground; so let s = 324 and solve for t.
--------
-16t^2+144t = 324
16t^2-144t+324 = 0
4(4t^2-36t+81) = 0
(2t-9)^2 = 0
t = 9/2 or 4.5 seconds
==============
Cheers,
Stan H.